OpenStudy (anonymous):

In the first set of problems. Units and Dimensional Analysis Challenge Problem Solutions - Problem 1: How did they find b(h/g)^1/2 ?

OpenStudy (anonymous):

If you follow the solution closely, the X and W are the exponents of g and h, respectively. Solving for these variables gave the values X = -1/2 and W = 1/2. As for b, it is a constant and therefore unitless so you have to add it in. This is why you get \[b \times (h ^{W}) \times (g^{X}) = b \times h ^{1/2} \times g^{-1/2}\] Hence, you get \[b(h/g)^{1/2}\] I hope this helps :)